# Arbitrage-free catastrophe reinsurance valuation for compound dynamic contagion claims

### v1

@@ 3.2 Special Case: Exponential/Gamma Distributions for Jump Sizes
$a$ $\rightarrow\ \theta\Bigl{(}1+\frac{\nu}{\gamma}\Bigr{)}^{-\eta}\frac{\beta}{\beta-B(t)}a;$
$\rho$ $\rightarrow\ \psi\frac{\alpha}{\alpha-B(t)}~{}\rho;$
$J\sim\mathsf{Gamma}(\gamma,\eta)$ $\rightarrow\ \mathsf{Gamma}(\gamma+\nu,\eta);$
$H\sim\mathsf{Exp}(\alpha)$ $\rightarrow\ \mathsf{Exp}(\lambda_{h});$
$G\sim\mathsf{Exp}(\beta)$ $\rightarrow\ \mathsf{Exp}(\lambda_{g});$
@@
Similarly, for gamma distribution $J\sim\mathsf{Gamma}(\gamma,\eta)$, we have the transform to $\mathsf{Gamma}(\gamma,\eta+\nu)$, since
@@
In fact, the original distributional structure of the underlying risk process via this change of measure is preserved. We may also need the stationary condition under the new measure, i.e.,
$\delta~{}~{}>~{}~{}\frac{\theta\frac{\gamma}{\gamma+\nu}\frac{\beta}{\beta-B(t)}}{\beta-B(t)}.$
So, we have to set $\delta$ relative large together with the condition $\delta>\theta~{}\hat{j}(\nu)\frac{1}{\beta}$.
@@
where $b>0,{B^{+}=\beta-\frac{1}{\delta}\theta~{}j(\nu)>0}$. Note that the function $\mathcal{G}(\cdot)$ in Eq. 3.13 shall be derived fully analytically. However, its inverse $\mathcal{G}^{-1}(\cdot)$ has to be solved numerically.
@@
$\mathbb{E}^{\ast}[\lambda_{t}\mid\lambda_{0}]$ $=\ \lambda_{0}\mathrm{e}^{-\int_{0}^{t}I(s)\mathop{}\!\mathrm{d}s}+\mathrm{e}^{-\int_{0}^{t}I(s)\mathop{}\!\mathrm{d}s}\theta\left(\frac{\gamma}{\gamma+\nu}\right)^{\eta}$
×$\quad\times\int_{0}^{t}\mathrm{e}^{\int_{0}^{s}I(u)\mathop{}\!\mathrm{d}u}\left(\frac{\beta}{\beta-B(s)}\right)\left\{\psi\alpha\rho\left(\frac{1}{\alpha-B(s)}\right)^{2}+a\delta\right\}\mathop{}\!\mathrm{d}s.$
$\mathbb{E}^{\ast}[N_{t}\mid\lambda_{0}]=\int_{0}^{t}\mathbb{E}^{\ast}[\lambda_{s}\mid\lambda_{0}]\mathop{}\!\mathrm{d}s$
$\mathbb{E}^{\ast}[C_{t}\mid\lambda_{0}]=\mu_{\widetilde{J}}\int_{0}^{t}\mathbb{E}^{\ast}[\lambda_{s}\mid\lambda_{0}]\mathop{}\!\mathrm{d}s$
@@ 4 Numerical Results
In this section, we provide arbitrage-free catastrophe stop-loss reinsurance premiums via the Monte Carlo simulation method. We also examine the sensitivity analyses for the Esscher parameters and the retention level.
@@ 4.1 Pricing of a stop-loss reinsurance contract
The total loss excess over $L$, which is a retention limit, up to time $t$ is $(C_{t}-L)^{+}$. Then the stop-loss reinsurance gross premium at time 0 (i.e. under $\mathbb{P}^{*}$) is
$\mathbb{E}^{\ast}[(C_{t}-L)^{+}].$ (4.1)
Its Monte Carlo estimate will be denoted by $\widehat{\mathbb{E}}^{\ast}[(C_{t}-L)^{+}]$, and the corresponding estimate under $\mathbb{P}$ will be denoted by $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$. For this section, we consider
@@
For each combination of parameters, 10,000 crude Monte Carlo (CMC) samples of $C_{t}$ under $\mathbb{P}^{*}$ have been used to estimate Eq. 4.1 in Table 1. The corresponding values under $\mathbb{P}$ are also provided in the same table.
Figure 1: A collection of 25 sample paths of $C_{t}$ and the corresponding $\lambda_{t}$ for the compound dynamic contagion process under the original measure $\mathbb{P}$ (the left column) and the tilted measure $\mathbb{P}^{\ast}$ (the right column) given the constants outlined above.
$L$ $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$ $\widehat{\mathbb{E}}^{\ast}[(C_{t}-L)^{+}]$
0.00 14.041136 38.152252
25.00 2.632637 19.153988
38.15 1.015409 12.894507
50.00 0.424692 8.980795
75.00 0.070956 4.113723
100.00 0.007726 1.859676
Table 1: The stop loss CMC estimates $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$ under the original measure and $\widehat{\mathbb{E}}^{\ast}[(C_{t}-L)^{+}]$ under the Esscher measure for different $A$ retention levels.
Table 1 shows that arbitrage-free catastrophe stop-loss reinsurance (i.e. gross) premiums are significantly higher than corresponding net premiums (see Fig. 1). By different combinations of parameters, the gross premiums could be lowered or elevated. It is not the purpose of this paper to decide which is the appropriate one to use. The insurance companies’ attitude towards risk determines which equivalent martingale probability measure should be used. The attractive thing about the Esscher transform is that it provides us with at least one equivalent martingale probability measure in incomplete market situations.
@@ 4.2 Sensitivity Analysis for the Esscher parameters
To examine the sensitivity analyses for the Esscher parameters, we focus on the mean, $\widehat{\mathbb{E}}^{\ast}[C_{t}]$ and the stop-loss retention $L=25$ for the above combination of parameters. Hence we have:
$\widehat{\mathbb{E}}^{\ast}[C_{t}]=38.152252\quad\text{ and }\quad\widehat{\mathbb{E}}^{\ast}[(C_{t}-25)^{+}]=19.153988\,,$
where $\widehat{\mathbb{E}}^{\ast}[C_{t}]$ itself can be used for arbitrage-free catastrophe insurance premiums.
The following tables show the way in which these values change when the Esscher parameters are adjusted individually, with all other parameters held constant (except $\phi$ which updates to always satisfy $\phi=-(\theta\hat{j}(\nu)-1)$).
@@ 4.2.1 Changing $\theta$
$\theta$ $\mathbb{E}^{\ast}[C_{t}]$ $\widehat{\mathbb{E}}^{\ast}[C_{t}]$ $\widehat{\mathbb{E}}^{\ast}[(C_{t}-25)^{+}]$
1.00 28.136544 28.55 $\pm$ 0.52 11.52 $\pm$ 0.41
1.25 37.756014 38.15 $\pm$ 0.68 19.15 $\pm$ 0.58
1.50 49.413171 49.82 $\pm$ 0.87 29.32 $\pm$ 0.80
1.75 63.671486 64.39 $\pm$ 1.11 42.71 $\pm$ 1.05
Table 2: The expected loss $\mathbb{E}^{\ast}[C_{t}]$, the CMC loss $\widehat{\mathbb{E}}^{\ast}[C_{t}]$ and stop-loss $\widehat{\mathbb{E}}^{\ast}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\theta$.
@@ 4.2.2 Changing $\psi$
$\psi$ $\mathbb{E}^{\ast}[C_{t}]$ $\widehat{\mathbb{E}}^{\ast}[C_{t}]$ $\widehat{\mathbb{E}}^{\ast}[(C_{t}-25)^{+}]$
1.00 34.774768 34.86 $\pm$ 0.65 16.72 $\pm$ 0.55
1.25 37.756014 38.15 $\pm$ 0.68 19.15 $\pm$ 0.58
1.50 40.736683 41.31 $\pm$ 0.71 21.61 $\pm$ 0.62
1.75 43.717393 44.86 $\pm$ 0.74 24.47 $\pm$ 0.66
Table 3: The expected loss $\mathbb{E}^{\ast}[C_{t}]$, the CMC loss $\widehat{\mathbb{E}}^{\ast}[C_{t}]$ and stop-loss $\widehat{\mathbb{E}}^{\ast}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\psi$.
@@ 4.2.3 Changing $\nu$
$\nu$ $\mathbb{E}^{\ast}[C_{t}]$ $\widehat{\mathbb{E}}^{\ast}[C_{t}]$ $\widehat{\mathbb{E}}^{\ast}[(C_{t}-25)^{+}]$
-0.01 22.321852 22.67 $\pm$ 0.42 7.26 $\pm$ 0.30
-0.05 37.756014 38.15 $\pm$ 0.68 19.15 $\pm$ 0.58
-0.08 62.101958 62.48 $\pm$ 1.08 41.04 $\pm$ 1.02
-0.10 93.448304 95.16 $\pm$ 1.67 72.32 $\pm$ 1.62
Table 4: The expected loss $\mathbb{E}^{\ast}[C_{t}]$, the CMC loss $\widehat{\mathbb{E}}^{\ast}[C_{t}]$ and stop-loss $\widehat{\mathbb{E}}^{\ast}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\nu$.

### v2

@@ 3.2 Infinitesimal generator under the equivalent martingale measure
From an economic perspective, the Esscher parameters $(\theta,\psi,\nu)$ provide a parsimonious representation of reinsurer risk preferences under hard market conditions. In particular, the transformed intensity process
$\tilde{\lambda}_{t}:=\theta\hat{j}(\nu)\hat{g}(-B(t))\,\lambda_{t}$
introduces a time-dependent loading on claim arrivals, reflecting heightened sensitivity to both claim frequency and loss severity during periods of elevated risk. The parameter $\theta$ acts as a frequency loading on the baseline arrival intensity, while the factor $\hat{j}(\nu)\hat{g}(-B(t))$ captures the interaction between severity aversion and contagion-driven dynamics. Similarly, the parameter $\psi$ scales the arrival rate of externally driven shocks, representing increased concern about systemic events. The Esscher parameter $\nu<0$ induces a severity-based loading by tilting the claim size distribution toward larger losses, thereby increasing the contribution of catastrophic events to the premium.
In a hard reinsurance market, these choices of the parameters are consistent with observed outcomes such as higher premiums and tighter underwriting, resulting in a conservative valuation. On the other hand, reinsurers may also limit coverage in high-risk regions by transferring part of the risk whenever alternative capacity is available. In this sense, the Esscher parameters $(\theta,\psi,\nu)$ may serve as reduced-form measures incorporating market risk aversion and capital scarcity into catastrophe reinsurance pricing.
@@ 3.3 Special Case: Exponential/Gamma Distributions for Jump Sizes
$a\$$\rightarrow\ \theta\Bigl(1+\frac{\nu}{\gamma}\Bigr)^{-\eta}\frac{\beta}{\beta-B(t)}a;$
$\rho\$$\rightarrow\ \psi\frac{\alpha}{\alpha-B(t)}~\rho;$
$J\sim\mathsf{Gamma}(\gamma,\eta)\$$\rightarrow\ \mathsf{Gamma}(\gamma+\nu,\eta);$
$H\sim\mathsf{Exp}(\alpha)\$$\rightarrow\ \mathsf{Exp}(\lambda_{h});$
$G\sim\mathsf{Exp}(\beta)\$$\rightarrow\ \mathsf{Exp}(\lambda_{g});$
@@
Similarly, for gamma distribution $J\sim\mathsf{Gamma}(\gamma,\eta)$, we have the transform to $\mathsf{Gamma}(\gamma+\nu,\eta)$, since
@@
In fact, the original distributional structure of the underlying risk process via this change of measure is preserved. We may also need the stationary condition for a CDCP under the new measure, i.e.,
$\delta~~>~~\frac{1}{\lambda_{g}}~~=~~\theta\left(\frac{\gamma}{\gamma+\nu}\right)^{\eta}\frac{\beta}{(\beta-B(t))^{2}},$
where $\delta$ and $\frac{1}{\lambda_{g}}$ are the decay rate and the mean of exponentially distributed self-excited jump sizes under the new measure for a CDCP, respectively, according to Corollary 3.3. So, for implementations, we have to set the decay rate $\delta$ relatively large together with the condition $\delta>\theta~\hat{j}(\nu)\mu_{G}=\theta~\hat{j}(\nu)\frac{1}{\beta}$ required in Theorem 3.2 for the existence and uniqueness of the solutions of the ODEs to hold. This constraint on $\delta$ is required in our model. In other words, the model may not be applicable if this constraint on $\delta$ does not hold. This may represent a limitation of the model.
@@
where $b>0,B^{+}=\beta-\frac{1}{\delta}\theta~\hat{j}(\nu)>0$. Note that the function $\mathcal{G}(\cdot)$ in Eq. 3.17 shall be derived fully analytically. However, its inverse $\mathcal{G}^{-1}(\cdot)$ has to be solved numerically.
@@
$\tilde{\mathbb{E}}[\lambda_{t}\mid\lambda_{0}]$ $=\ \lambda_{0}\mathrm{e}^{-\int_{0}^{t}I(s)\mathop{}\!\mathrm{d}s}+\mathrm{e}^{-\int_{0}^{t}I(s)\mathop{}\!\mathrm{d}s}\theta\left(\frac{\gamma}{\gamma+\nu}\right)^{\eta}$
$\quad\times\int_{0}^{t}\mathrm{e}^{\int_{0}^{s}I(u)\mathop{}\!\mathrm{d}u}\left(\frac{\beta}{\beta-B(s)}\right)\left\{\psi\alpha\rho\left(\frac{1}{\alpha-B(s)}\right)^{2}+a\delta\right\}\mathop{}\!\mathrm{d}s.$
$\tilde{\mathbb{E}}[N_{t}\mid\lambda_{0}]=\int_{0}^{t}\tilde{\mathbb{E}}[\lambda_{s}\mid\lambda_{0}]\mathop{}\!\mathrm{d}s$
$\tilde{\mathbb{E}}[C_{t}\mid\lambda_{0}]=\mu_{\widetilde{J}}\int_{0}^{t}\tilde{\mathbb{E}}[\lambda_{s}\mid\lambda_{0}]\mathop{}\!\mathrm{d}s$
@@
For illustrative purposes, we have used a Gamma distribution for claim sizes, which yields a closed-form expression under the Esscher transform, $\tilde{J}\sim\mathsf{Gamma}(\gamma+\nu,\eta)$. The Gamma family is a standard choice in actuarial modeling of claim severities (Bühlmann, 1970, Page 7) and offers analytical tractability that facilitates clear interpretation of the pricing mechanism. While heavier-tailed alternatives, such as the Pareto distribution,
$j(\xi)=\frac{\Gamma(\omega+k)\,\zeta^{\omega}\,\xi^{k-1}}{\Gamma(\omega)\,\Gamma(k)\,(\zeta+\xi)^{\omega+k}},\qquad\omega>0,\ \zeta>0,\ k>0,$
are commonly employed to model large insurance losses (Albrecher et al., 2017, Chapters 3–4), their form is not preserved under the Esscher transform and instead acquires an exponential tilt. Nevertheless, the corresponding Laplace transform $\hat{j}(\nu)$ exists under suitable parameter conditions and can be evaluated numerically. Heavier-tailed claim size distributions typically lead to higher premium levels, reflecting increased tail risk. In this sense, the Gamma specification, including the exponential distribution as a special case, serves as a benchmark for assessing the impact of tail heaviness on catastrophe reinsurance pricing. The results presented here therefore illustrate the qualitative behavior of the model, while extensions to Pareto or other heavy-tailed distributions may be explored in future work via numerical sensitivity analysis.
@@ 4 Numerical results
In this section, we provide arbitrage-free catastrophe stop-loss reinsurance premiums via the Monte Carlo simulation method. We also examine the sensitivity analyses for the retention level, the Esscher parameters and the intensity parameters.
@@ 4.1 Pricing of a stop-loss reinsurance contract
The total loss excess over $L$, which is a retention limit, up to time $t$ is $(C_{t}-L)^{+}$. Then the stop-loss reinsurance gross premium at time 0 (i.e. under $\tilde{\mathbb{P}}$) is
$\tilde{\mathbb{E}}[(C_{t}-L)^{+}].$ (4.1)
Its Monte Carlo estimate will be denoted by $\widehat{\tilde{\mathbb{E}}}[(C_{t}-L)^{+}]$, and the corresponding estimate under $\mathbb{P}$ will be denoted by $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$. For this section, we consider
@@
The parameter values adopted in the numerical examples are chosen for illustrative purposes. They are used to demonstrate the qualitative behaviour of the model, as well as the impacts of contagion and risk loading on reinsurance premiums, rather than to provide an empirical calibration to a specific portfolio. Similar parameter ranges are commonly adopted in the dynamic contagion and self-exciting point process literature to study clustering effects and catastrophic risk (Dassios and Zhao (2011); Jang and Oh (2021, 2025)).
In practice, model parameters may be estimated from historical catastrophe or large-loss data using standard inference techniques for point processes with latent or stochastic intensities, including likelihood-based methods, moment-based estimators, and filtering approaches. General statistical frameworks for point process inference are discussed in Karr (1991) and Daley and Vere-Jones (2008), while modern likelihood-based methodologies for self-exciting and Hawkes-type processes are presented in Laub et al. (2022). Related filtering approaches in insurance risk models are studied in Dassios and Jang (2005) for shot-noise Cox processes, and similar ideas may be adapted to the dynamic contagion setting considered here.
While maximum likelihood estimators for point process parameters and some of their statistical properties have been studied, the estimation of shot-noise Cox processes, and in particular the analysis of their asymptotic properties, has received comparatively less attention. These challenges are further compounded when time-varying parameters are present, suggesting promising opportunities for future research, particularly in the context of dynamic contagion models for insurance risk.
For each combination of parameters, $10^{5}$ crude Monte Carlo (CMC) samples of $C_{t}$ under $\tilde{\mathbb{P}}$ have been used to estimate Eq. 4.1 in Table 1. The corresponding values under $\mathbb{P}$ are also provided in the same table.
Figure 1: A collection of 25 sample paths of $C_{t}$ and the corresponding $\lambda_{t}$ for the compound dynamic contagion process under the original measure $\mathbb{P}$ (left column) and the tilted measure $\tilde{\mathbb{P}}$ (right column) given the constants outlined above. It indicates that paths under $\tilde{\mathbb{P}}$ exhibit higher intensities and larger claims, reflecting risk loading and increased emphasis on extreme losses.
$L$ $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-L)^{+}]$
0.00 13.867646 37.635418
25.00 2.556228 18.752166
37.64 1.011775 12.792098
50.00 0.406231 8.738347
75.00 0.062353 4.023879
100.00 0.010058 1.852134
Table 1: The stop-loss CMC estimates $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$ under the original measure and $\widehat{\tilde{\mathbb{E}}}[(C_{t}-L)^{+}]$ under the Esscher measure for different $L$ retention levels.
Table 1 shows that arbitrage-free catastrophe stop-loss reinsurance (i.e. gross) premiums are significantly higher than the corresponding net premiums (see Fig. 1). Depending on the choice of parameters, the gross premiums could be either lowered or elevated. It is not the purpose of this paper to determine which premium is the most appropriate; rather, the insurance companies’ risk preferences dictate which equivalent martingale probability measure should be adopted.
The simplest way to obtain a gross premium is through the use of the parameter $\theta$, a well-known concept in insurance where a security loading is added to the expected value under the original probability measure $\mathbb{P}$. An appealing feature of the Esscher transform is that it guarantees the existence of at least one equivalent martingale probability measure in incomplete markets.
@@ 4.2 Sensitivity analysis for the Esscher parameters
To examine the sensitivity analyses for the Esscher parameters, we focus on the mean, $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ and the stop-loss retention $L=25$ for the above combination of parameters. Hence we have:
$\widehat{\tilde{\mathbb{E}}}[C_{t}]=38.152252\quad\text{ and }\quad\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]=19.153988\,$,
where $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ itself can be used for arbitrage-free catastrophe insurance premiums.
The following tables show the way in which these values change when the Esscher parameters are adjusted individually, with all other parameters held constant (except $\phi$ which is updated so that it always satisfies $\phi=-(\theta\hat{j}(\nu)-1)$).
@@ 4.2.1 Changing $\theta$
$\theta$ $\tilde{\mathbb{E}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$
1.00 28.137195 28.12 ± 0.16 11.27 ± 0.13
1.25 37.757126 37.64 ± 0.21 18.75 ± 0.18
1.50 49.413007 49.25 ± 0.27 28.80 ± 0.25
1.75 63.671363 63.58 ± 0.35 41.95 ± 0.33
Table 2: The expected loss $\tilde{\mathbb{E}}[C_{t}]$, the CMC loss $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ and stop-loss $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\theta$.
@@ 4.2.2 Changing $\psi$
$\psi$ $\tilde{\mathbb{E}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$
1.00 34.775140 34.84 ± 0.21 16.69 ± 0.17
1.25 37.757126 37.64 ± 0.21 18.75 ± 0.18
1.50 40.737259 40.67 ± 0.22 21.13 ± 0.19
1.75 43.718045 43.83 ± 0.23 23.66 ± 0.21
Table 3: The expected loss $\tilde{\mathbb{E}}[C_{t}]$, the CMC loss $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ and stop-loss $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\psi$.
@@ 4.2.3 Changing $\nu$
$\nu$ $\tilde{\mathbb{E}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$
-0.01 22.322487 22.39 ± 0.13 7.13 ± 0.09
-0.05 37.757126 37.64 ± 0.21 18.75 ± 0.18
-0.08 62.101734 61.90 ± 0.34 40.47 ± 0.32
-0.10 93.448189 93.15 ± 0.52 70.32 ± 0.51
Table 4: The expected loss $\tilde{\mathbb{E}}[C_{t}]$, the CMC loss $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ and stop-loss $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\nu$.
@@
4.3 Sensitivity analysis for the intensity parameters
We next examine sensitivity with respect to the intensity parameters governing the dynamics of the claim arrival process. Using the same setup (i.e. the same baseline parameter values) as in the previous subsection, we focus on the mean $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ and the stop-loss premium with retention $L=25$.
The following tables illustrate how these quantities change when key intensity parameters are varied individually, with all other parameters held constant (except $\phi$, which is updated to satisfy $\phi=-(\theta\hat{j}(\nu)-1)$).
4.3.1 Changing $\delta$
$\delta$ $\tilde{\mathbb{E}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$
2.00 43.634528 43.60 ± 0.26 24.23 ± 0.23
3.00 37.757126 37.64 ± 0.21 18.75 ± 0.18
5.00 32.620608 32.69 ± 0.18 14.26 ± 0.14
7.00 58.289065 58.21 ± 0.37 37.36 ± 0.35
Table 5: The expected loss $\tilde{\mathbb{E}}[C_{t}]$, the CMC loss $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ and stop-loss $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\delta$.
4.3.2 Changing $\alpha$
$\alpha$ $\tilde{\mathbb{E}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$
1.00 53.134574 53.19 ± 0.27 32.01 ± 0.25
2.00 37.757126 37.64 ± 0.21 18.75 ± 0.18
3.00 32.736742 32.74 ± 0.20 15.03 ± 0.16
4.00 30.246361 30.30 ± 0.19 13.28 ± 0.15
Table 6: The expected loss $\tilde{\mathbb{E}}[C_{t}]$, the CMC loss $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ and stop-loss $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\alpha$.
4.3.3 Changing $\beta$
$\beta$ $\tilde{\mathbb{E}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$
1.00 37.757126 37.64 ± 0.21 18.75 ± 0.18
1.50 31.529562 31.50 ± 0.16 13.08 ± 0.13
2.00 29.013963 29.04 ± 0.15 10.90 ± 0.11
2.50 27.659677 27.73 ± 0.14 9.77 ± 0.10
Table 7: The expected loss $\tilde{\mathbb{E}}[C_{t}]$, the CMC loss $\widehat{\tilde{\mathbb{E}}}[C_{t}]$ and stop-loss $\widehat{\tilde{\mathbb{E}}}[(C_{t}-25)^{+}]$ estimates (with 95% confidence intervals), for different values of the tilting parameter $\beta$.
Table 5 shows the complicated relationship that $\delta$ has on the expected losses. Changing $\delta$ changes the solution $B(t)$, and that changes nearly all of the Esscher-transformed parameters (except for claim sizes), making the impacts non-monotonic. Tables 6 and 7 present the reduction in them as the external-exciting and self-exciting jump size parameters $\alpha$ and $\beta$ increase. An increase in these jump size parameters results in a decrease in the exponential average.
4.4 Comparison with Generalised Hawkes and Shot-noise Cox processes
If there are no externally excited jumps in Section 2.2.1 (i.e., $\rho=0$) $C_{t}$ becomes generalised compound Hawkes process. If $a=0$ and there are no self-excited jumps in Section 2.2.1, $C_{t}$ becomes compound Cox process with shot-noise Poisson intensity.
The following two tables show the corresponding stop-loss CMC estimates in Table 1 (DCP case) for the generalised Hawkes case and the Cox case, respectively.
4.4.1 Generalised Hawkes process
If we simulate using the same parameters as above, except setting $\rho=0$, then the underlying point process becomes a generalised Hawkes process.
$L$ $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-L)^{+}]$
0.00 9.643789 22.869498
25.00 1.251784 8.592431
37.64 0.441185 5.278698
50.00 0.160856 3.318908
75.00 0.019958 1.351518
100.00 0.002533 0.573990
Table 8: The stop-loss CMC estimates $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$ under the original measure and $\widehat{\tilde{\mathbb{E}}}[(C_{t}-L)^{+}]$ under the Esscher measure for different $L$ retention levels. The point process driving $C_{t}$ is a generalised Hawkes process.
4.4.2 Shot-noise Cox process
If we simulate using the same parameters as above, except setting $a=0$ and $\beta=\infty$ (i.e., $Y_{j}\equiv 0$ for all $j$), then the underlying point process becomes a shot-noise Cox process.
$L$ $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$ $\widehat{\tilde{\mathbb{E}}}[(C_{t}-L)^{+}]$
0.00 5.804331 12.216566
25.00 0.231064 1.840665
37.64 0.035712 0.608449
50.00 0.004792 0.192236
75.00 0.000030 0.017000
100.00 0.000000 0.001662
Table 9: The stop-loss CMC estimates $\widehat{\mathbb{E}}[(C_{t}-L)^{+}]$ under the original measure and $\widehat{\tilde{\mathbb{E}}}[(C_{t}-L)^{+}]$ under the Esscher measure for different $L$ retention levels. The point process driving $C_{t}$ is a shot-noise Cox process.
Compared to Tables 8 and 9, Table 1 shows that gross insurance and reinsurance premium estimates significantly increase when changing $C_{t}$ from the Cox process with shot-noise Poisson intensity to a generalised Hawkes process due to self-excited jumps, and to the dynamic contagion process due to both externally excited jumps and self-excited jumps.